OpenStudy (anonymous):

Use a calculator to determine the indicated acute angle to the nearest tenth of a degree. sin A = 0.6789 how do you plug it into the calculator?

OpenStudy (anonymous):

\[\sin A = 0.6789 \] or \[A = \sin ^{-1}0.6789 \] get that?there must be a \[\sin ^{-1}\] key on your calculator

OpenStudy (anonymous):

ohh okay i found the button thank you so much!

OpenStudy (wolf1728):

Incidentally if the sine of an angle is .6789 to find the angle you look up the arc sine of .6789